# TokenSelect: Efficient Long-Context Inference and Length Extrapolation for LLMs via Dynamic Token-Level KV Cache Selection

### v1

@@ 3.2. Selective Sparse Attention
Figure 2. Motivations for token-level selection. (a) Visualization of attention scores sparsity. (b) Attention scores and critical token recalled by 1K token budget. (c) The $L_{1}$ norm of attention logits in each attention head. Visualizations are based on Qwen-2-7B-Instruct on the GovReport dataset of LongBench.
@@
(a) Consecutive query is similar and the similarity distribution is consistent across datasets.
(b) Selection overlaps when query is similar.
@@
Figure 3. Observations on similarity of consecutive queries. (a) Cosine similarity distribution between consecutive queries. (b) The token selection overlap rate ($\frac{|\mathcal{I}_{i}\cap\mathcal{I}_{i+1}|}{|\mathcal{I}_{i+1}|}$) with respect to consecutive query similarity.
Definition 1 (Selective Sparse Attention Problem, informal). For current input of length $C$ ($C=1$ in the Decode Stage) and KV Cache of length $N$, assuming there are $H$ attention heads with a head size of $d_{h}$, let $\mathbf{O}$ be the output of the SDPA:
(4) $\mathbf{O}=\left[\text{softmax}\left(\frac{\mathbf{Q}^{h}\cdot{\left[\mathbf{K}^{h}_{\text{cache}},\ \mathbf{K}^{h}_{\text{current}}\right]}^{\top}}{\sqrt{d}}\right)\cdot[\mathbf{V}^{h}_{\text{cache}},\ \mathbf{V}^{h}_{\text{current}}]\right]_{h\in\{1,\cdots,H\}},$
where $\mathbf{Q}^{h},\mathbf{K}^{h}_{\text{current}},\mathbf{V}^{h}_{\text{current}}\in\mathbb{R}^{C\times d_{h}}$ is the Query, Key, Value matrices of current input for head $h$ and $\mathbf{K}^{h}_{\text{cache}},\mathbf{V}^{h}_{\text{cache}}\in\mathbb{R}^{N\times d_{h}}$ represent the KV Cache. Let $\hat{\mathbf{O}}$ be the output of the Selective Sparse Attention:
(5) $\hat{\mathbf{O}}=\left[\text{softmax}\left(\frac{\mathbf{Q}^{h}\cdot{\left[\mathbf{K}^{h}_{\text{select}},\ \mathbf{K}^{h}_{\text{current}}\right]}^{\top}}{\sqrt{d}}\right)\cdot[\mathbf{V}^{h}_{\text{select}},\ \mathbf{V}^{h}_{\text{current}}]\right]_{h\in\{1,\cdots,H\}},$
(6) $\mathcal{S}\left(\mathbf{Q},\ \mathbf{K}_{\text{cache}}\right)=\mathcal{I},\;\text{where}\;\mathcal{I}\in\mathcal{P}(\{1,\cdots,N\}),$
(7) $\min_{S}\left\|\mathbf{O}-\hat{\mathbf{O}}\right\|_{2}^{2}.$
Existing works on long-context inference (Wang et al., 2019; Zaheer et al., 2020; Xiao et al., 2024; Han et al., 2024; Zhang et al., 2024; Oren et al., 2024; Li et al., 2024; Xiao et al., 2024; Tang et al., 2024; Jiang et al., 2024) can be categorized under the Selective Sparse Attention Problem, with variations primarily in the design of the selection function $\mathcal{S}$.  (Wang et al., 2019; Zaheer et al., 2020; Xiao et al., 2024) have developed input-independent selection functions $\mathcal{S}()$, while  (Zhang et al., 2024; Oren et al., 2024; Li et al., 2024) propose query-independent functions $\mathcal{S}(\mathbf{K}_{\text{cache}})$ for improved performance. Current state-of-the-art methods (Xiao et al., 2024; Tang et al., 2024; Jiang et al., 2024) utilize query-aware selection functions $\mathcal{S}(\mathbf{Q},\mathbf{K}_{\text{cache}})$. However, these approaches typically operate at a block-level, which limits their effectiveness and overall performance.
4. Motivations and Observations
Attention is Sparse, Non-contiguous and Head-Distinctive
@@ 4. Motivations and Observations
Previous works (Wang et al., 2019; Zaheer et al., 2020; Xiao et al., 2024; Han et al., 2024; Zhang et al., 2024; Oren et al., 2024; Li et al., 2024; Xiao et al., 2024; Tang et al., 2024; Jiang et al., 2024) have demonstrated the sparsity of attention scores in LLMs, particularly when processing long texts. Recent approaches (Xiao et al., 2024; Jiang et al., 2024; Tang et al., 2024) partition the KV Cache into non-overlapping blocks, estimating block criticality for sparse attention calculations. These methods assume that tokens with higher attention scores tend to be contiguous. However, our further observations reveal that this assumption does not always hold true in practice. As illustrated in Fig. 2a, attention scores are sparsely distributed at the token-level, with critical tokens not necessarily contiguous. This non-contiguity leads to significant omissions in block-level token selection. Fig. 2b demonstrates that finer selection granularity improves recall of critical tokens, motivating us to perform token-level selection. For token-level selection, an intuitive approach would be to directly select the top-$k$ tokens with the highest attention logits. However, observation in Fig. 2c reveals considerable disparity in the $L_{1}$ norm of attention logits across attention heads. As a result, the selection result tends to be dominated by a few heads with disproportionately large attention logits, driving us to design a more robust selection function that maintains the independence of heads.
Consecutive Queries are similar
@@
As sparsity of attention is dynamic (Zhang et al., 2024; Xiao et al., 2024; Tang et al., 2024; Jiang et al., 2024), token selection should be performed for every Query, which inevitably increases the computational overhead of selective sparse attention. Fortunately, we observe that consecutive Queries exhibit high similarity, as shown in Fig. 3a. Intuitively, when two consecutive Queries are highly similar, their dot products with the Keys will also be similar, leading to substantial overlap in the token selection results. Due to space constraints, we provide an informal lemma about this below. The formal version and corresponding proof can be found in the Appendix A.
Figure 4. The illustration of TokenSelect, which involves calculating per-head criticality using the Paged Dot Product Kernel, performing head soft vote to get selected indices, and executing selective sparse attention via the Paged Attention Kernel.
@@
Lemma 1 (Informal). Consider Queries $\mathbf{Q}_{1},\mathbf{Q}_{2}\in\mathbb{R}^{1\times d}$ that are consecutive and a Key set $\{\mathbf{K}_{i}\}_{i=1}^{N}$. Let $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ be the sets of indices of the top-$k$ Keys selected by dot product for $\mathbf{Q}_{1}$ and $\mathbf{Q}_{2}$ respectively. If $cos(\mathbf{Q}_{1},\mathbf{Q}_{2})>\epsilon$, where $\epsilon$ is a threshold, then $\mathcal{I}_{1}=\mathcal{I}_{2}$.
@@
Fig. 3b illustrates this lemma experimentally. It can be seen that the overlap rate of token selection tends to increase with query similarity. This key insight motivates us to reuse selection results for similar queries, improving computational efficiency. Moreover, the similarity distribution of consecutive Queries remains consistent across different tasks, as demonstrated in Fig. 3a, allowing us to apply a global similarity threshold across all scenarios.
5. Designs of TokenSelect
@@
5.1. Selection Function
The simplest selection function is to determine the criticality of the tokens through the dot product of $\mathbf{Q}$ and $\mathbf{K}_{\text{cache}}$, then select the top-$k$ critical ones as $\mathbf{K}_{\text{select}},\mathbf{V}_{\text{select}}$. The selected indices $\mathcal{I}$ are calculated as follow:
(8) $\mathcal{I}_{\text{topk}}=\operatorname{TopK}\left(\mathbf{Q}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right).$
@@ 5.1. Selection Function
(9) $\mathcal{I}_{\text{head-vote}}=\operatorname{TopK}\left(\sum_{h=1}^{H}\mathbb{I}\left(i\in\operatorname{TopK}\left(\mathbf{Q}^{h}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right)\right)\right),$
where $\mathbb{I}$ is the indicator function. Unfortunately, despite better performance, this method relies on scatter_add and multiple topk operations, resulting in low efficiency on GPUs. Additionally, the 0/1 voting ignores the relative importance of tokens for each head. Therefore, we propose a head soft vote approach that offers better performance and efficiency. Specifically, we first calculate the per-head criticality, then normalize through softmax, and sum the results for all heads. The formalization is as follows:
(10) $\mathcal{I}_{\text{head-soft-vote}}=\operatorname{TopK}\left(\sum_{h=1}^{H}\operatorname{softmax}\left(\mathbf{Q}^{h}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right)\right).$
5.2. Optimizing Selection Frequency
Although the aforementioned selection function can reduce the complexity of attention from $O(N^{2})$ to $O(k^{2})$ ($k\ll N$), while maintaining performance, the execution time of the selection function itself still affects the latency of LLMs inference. To further accelerate long-context inference, based on our observations of the similarity of consecutive queries, we design optimization strategies for both the Prefill Stage and the Decode Stage to reduce the selection frequency while ensuring its effectiveness.
In the Prefill Stage, $\mathbf{Q}_{\text{prefill}}\in\mathbb{R}^{n_{\text{in}}\times d}$ is inputed. In long-context scenarios, the number of tokens in the user’s input sequence $n_{\text{in}}$ may reach up to 1M, making it impractical to perform selection for each Query token. Considering the similarity of consecutive Queries, we use chunk-wise token selection, inputting $\frac{1}{c}\sum_{i=1}^{c}(\mathbf{Q}_{C})_{i}$ into the selection function, where $\mathbf{Q}_{C}\in\mathbb{R}^{c\times d}$ is the Query chunk and $c$ is the chunk size. This method helps maintain the compute-intensive nature of the Prefill Stage, preventing it from becoming memory bound.
Algorithm 1 Selection Cache Algorithm for the Decode Stage
0:  $\mathbf{Q}$: current Query,  $k$: number of selected tokens,      $\mathbf{C}_{Q}$: Query cache, $\mathbf{C}_{\mathcal{I}}$: selection cache,      $\mathcal{S}$: selection function (Eq.(10)), $\theta$: similarity threshold,      $\mathbf{f}$: first query flags (default True)
@@ 5.2. Optimizing Selection Frequency
In the Decode Stage, due to the auto-regressive characteristic of LLMs, we need to frequently perform selection for $\mathbf{Q}_{\text{decode}}\in\mathbb{R}^{1\times d}$, and this process cannot be executed chunk-wise like in the Prefill Stage. To reduce the frequency of token selection in the Decode Stage, as shown in Algorithm 1, we propose the Selection Cache. Consecutive similar Queries will hit the cache, thereby directly loading the cached selection results for the previous Query. The Selection Cache allows us to reduce decode latency while maintaining almost the same performance.
Table 1. Comparison of different methods with different origin models on InfiniteBench.
@@
5.3. Efficient Implementation
To ensure that our proposed TokenSelect can be used for real-world Web applications, efficient implementation is crucial. We first analyze the computation time breakdown of representative block-level selective sparse attention method, InfLLM (Xiao et al., 2024). From (1)(2)(3) in Fig. 5, we can observe that although selective sparse attention can significantly reduce the complexity of attention calculations, the actual computation time is still highly dependent on the implementation. The incompatibility with efficient attention implementations such as Flash Attention has resulted in methods requiring historical attention scores (Zhang et al., 2024; Oren et al., 2024; Li et al., 2024; Xiao et al., 2024) being difficult to be applied in real-world Web applications. Through the analysis of InfLLM’s Flash Attention-compatible version, we make several discoveries. The initial motivation for estimating token criticality at the block-level is to reduce the overhead of selection function (mainly considering dot product calculation). However, we find that dot product is not the primary performance bottleneck. Instead, a significant portion of the overhead comes from indexing the KV Cache using selected indices and making them contiguous in GPU memory, which frequently occurs during the updating of KV blocks and the concatenation of selected KV Cache. The extensive I/O required for this operation further exacerbates the memory-bound in LLMs inference. Based on this, we propose that Paged Attention is a more suitable implementation for selective sparse attention. Using Paged KV Cache management (with page size=1 for TokenSelect), we can reduce the I/O volume for selection results from the scale of all selected KV Caches $O(2kd)$ to the scale of their indices $O(k)$. However, by observing (4) in Fig. 5, we find that we encounter another bottleneck under Paged KV Cache management. Since logically contiguous KV Cache is not entirely contiguous in GPU memory, it also needs to be made contiguous before performing computational operations. To address this issue, we draw inspiration from the concept of Paged Attention and implement a Paged Dot Product Kernel using Triton (Tillet et al., 2019), which significantly improves the overall efficiency of TokenSelect.
Figure 5. Computation time breakdown for single chunk prefill step under different attention implementations (chunk size: 512, KV Cache length: 128K, attended tokens: 4K).
6. Experiments
In this section, we first introduce the experimental setup of this paper, and then reveal the performance and efficiency of our TokenSelect in long-context inference through experiments.
6.1. Experimental Settings
@@
To evaluate TokenSelect’s performance on long-context inference, we use the following datasets: (1) InfiniteBench (Zhang et al., 2024): The mainstream long-context benchmark consisting of multi-tasks. The average length of it exceeds 200K tokens. (2) RULER (Hsieh et al., 2024): A challenging long-context benchmark containing 13 different tasks, with subsets of varying lengths up to 128K tokens. (3) LongBench (Bai et al., 2024): Another mainstream long-context benchmark comprising 6 types of tasks. The 95% percentile for its lengths is 31K tokens. For each dataset, we use its recommended metrics, which are presented in the Appendix B.
@@
To demonstrate the state-of-the-art (SOTA) performance of TokenSelect, we include the following methods for comparison: (1) Original models: We select three mainstream open-source LLMs - Qwen2-7B-Instruct (Yang et al., 2024), Llama-3-8B-Instruct (Dubey et al., 2024), and Yi-1.5-6B-Chat (AI et al., 2024) - utilizing their original context lengths without any modifications. (2) NTK-Aware Scaled RoPE: A nonlinear RoPE interpolation method. (3) SelfExtend: A RoPE interpolation method that reuses the position ids across neighboring tokens. (4) StreamingLLM: The SOTA method for long-context inference with predefined sparse patterns. (5) InfLLM: The SOTA method for long-context inference and length extrapolation using a block-level selective sparse attention method. (6) MInference: The SOTA method for long-context prefilling acceleration, utilizing three sparse patterns including block-level sparse attention. It’s worth noting that since MInference doesn’t support length extrapolation, we use an alternative evaluation method, applying it to Llama-3-8B-Instruct-262k (Llama3 after long-text post-training). Additionally, we do not include another SOTA method, QUEST (Tang et al., 2024), as it does not support Grouped Query Attention (GQA).
Table 2. Comparison of different methods with different origin models on RULER.

### v2

@@ 4 Motivations and Observations
Figure 2: Motivations for token-level selection. (a) Visualization of attention scores sparsity. (b) Attention scores and critical token recalled by 1K token budget. (c) The $L_{1}$ norm of attention logits in each attention head.
@@
(a) Consecutive queries show consistent similarity patterns across datasets.
(b) Selection overlaps with similar queries.
@@
Figure 3: Observations on similarity of consecutive queries. (a) Cosine similarity distribution between consecutive queries. (b) The token selection overlap rate ($\frac{|\mathcal{I}_{i}\cap\mathcal{I}_{i+1}|}{|\mathcal{I}_{i+1}|}$) with respect to consecutive Query similarity.
Attention is Sparse, Non-contiguous and Head-Distinctive.
@@
Previous works on long-context inference have demonstrated the sparsity of attention scores in LLMs, particularly when processing long texts. Recent approaches Xiao et al. (2024a); Jiang et al. (2024); Tang et al. (2024b) partition the KV Cache into non-overlapping blocks, estimating block criticality for sparse attention calculations. These methods assume that tokens with higher attention scores tend to be contiguous. However, our further observations reveal that this assumption does not always hold true in practice. As illustrated in Fig. 2a, attention scores are sparsely distributed at the token-level, with critical tokens not necessarily contiguous. This non-contiguity leads to significant omissions in block-level token selection. Fig. 2b demonstrates that finer selection granularity improves recall of critical tokens, motivating us to perform token-level selection. For token-level selection, an intuitive approach would be to directly select the top-$k$ tokens with the highest attention logits. However, observation in Fig. 2c reveals considerable disparity in the $L_{1}$ norm of attention logits across attention heads. As a result, the selection result tends to be dominated by a few heads with disproportionately large attention logits, driving us to design a more robust selection function that maintains the independence of heads.
Consecutive Queries are Similar.
@@
As sparsity of attention is dynamic Jiang et al. (2024), token selection should be performed for every Query, which inevitably increases the computational overhead of selective sparse attention. Fortunately, we observe that consecutive Queries exhibit high similarity, as shown in Fig. 3a. Intuitively, when two consecutive Queries are highly similar, their dot products with the Keys will also be similar, leading to substantial overlap in the token selection results. Due to space constraints, we provide an informal lemma about this below. The formal version and corresponding proof can be found in the Appendix A.
@@
Lemma 1 (Informal). Consider Queries $\mathbf{Q}_{1},\mathbf{Q}_{2}\in\mathbb{R}^{1\times d}$ that are consecutive and a Key set $\{\mathbf{K}_{i}\}_{i=1}^{N}$. Let $\mathcal{I}_{1},$ and $\mathcal{I}_{2}$ be the sets of indices of the top-$k$ Keys selected by dot product for $\mathbf{Q}_{1},$ and $\mathbf{Q}_{2}$ respectively. If $cos(\mathbf{Q}_{1},\mathbf{Q}_{2})>\epsilon$, where $\epsilon$ is a threshold, then $\mathcal{I}_{1}=\mathcal{I}_{2}$.
@@
Fig. 3b illustrates this lemma experimentally. It shows that the overlap rate of token selection tends to increase with Query similarity. This key insight motivates us to reuse selection results for similar queries, improving computational efficiency. Moreover, the similarity distribution of consecutive Queries remains consistent across different tasks, as demonstrated in Fig. 3a, allowing us to apply a global similarity threshold across all scenarios.
Figure 4: The illustration of TokenSelect, which involves calculating per-head criticality using the Paged Dot Product Kernel, performing head soft vote to get selected indices, and executing selective sparse attention via the Paged Attention Kernel.
5 Designs of TokenSelect
@@
5.1 Selection Function
The simplest selection function is to determine the criticality of the tokens through the dot product of $\mathbf{Q}$ and $\mathbf{K}_{\text{cache}}$, then select the top-$k$ critical ones as $\mathbf{K}_{\text{select}},\mathbf{V}_{\text{select}}$. The selected indices $\mathcal{I}$ are calculated as follows:
$\mathcal{I}_{\text{topk}}=\operatorname{TopK}\left(\mathbf{Q}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right).$ (8)
@@ 5.1 Selection Function
$\mathcal{I}_{\text{head-vote}}=\operatorname{TopK}\left(\sum_{h=1}^{H}\mathbb{I}\left(i\in\operatorname{TopK}\left(\mathbf{Q}^{h}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right)\right)\right),$
where $\mathbb{I}$ is the indicator function. Unfortunately, despite better performance, this method relies on scatter_add and multiple topk operations, resulting in low efficiency on GPUs. Additionally, the 0/1 voting ignores the relative importance of tokens for each head. Therefore, we propose a head soft vote approach that offers better performance and efficiency. Specifically, we first calculate the per-head criticality, then normalize through softmax, and sum the results for all heads:
$\mathcal{I}_{\text{head-soft-vote}}=\operatorname{TopK}\left(\sum_{h=1}^{H}\sigma\left(\mathbf{Q}^{h}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right)\right).$
Algorithm 1 Selection Cache Algorithm
0:  $\mathbf{Q}$: current Query,  $k$: number of selected tokens,     $\mathbf{C}_{Q}$: Query cache, $\mathbf{C}_{\mathcal{I}}$: selection cache,      $\theta$: similarity threshold,      $\mathcal{S}$: selection function (Eq.(10)),      $\mathbf{f}$: first query flags (default True)
@@
5.2 Optimizing Selection Frequency
Although the aforementioned selection function can reduce the complexity of attention from $O(N^{2})$ to $O(k^{2}),k\ll N$, while maintaining performance, the execution time of the selection function itself still affects the latency of inference. To further accelerate long-context inference, based on our observations of the similarity of consecutive queries, we design optimization strategies for both the Prefill Stage and the Decode Stage to reduce the selection frequency while ensuring its effectiveness.
In the Prefill Stage, $\mathbf{Q}_{\text{prefill}}\in\mathbb{R}^{n_{\text{in}}\times d}$ is input. In long-context scenarios, the number of tokens in the user’s input sequence $n_{\text{in}}$ may reach up to 1M, making it impractical to perform selection for each Query token. Considering the similarity of consecutive Queries, we use chunk-wise token selection, inputting $\frac{1}{c}\sum_{i=1}^{c}(\mathbf{Q}_{C})_{i}$ into the selection function, where $\mathbf{Q}_{C}\in\mathbb{R}^{c\times d}$ is the Query chunk and $c$ is the chunk size. This method helps maintain the compute-intensive nature of the Prefill Stage, preventing it from becoming memory bound.
In the Decode Stage, due to the auto-regressive characteristic of LLMs, we need to frequently perform selection for $\mathbf{Q}_{\text{decode}}$, and this process cannot be executed chunk-wise like in the Prefill Stage. To reduce the frequency of token selection in the Decode Stage, we propose the Selection Cache. Consecutive similar Queries will hit the cache, thereby directly loading the cached selection results for the previous Query. The Selection Cache allows us to reduce decode latency while maintaining the performance. The formal formulation of the Selection Cache is detailed in Algorithm 1.
@@ 5.2 Optimizing Selection Frequency
Table 1: Comparison of different methods with different origin models on InfiniteBench.
5.3 Efficient Implementation
To ensure that our proposed TokenSelect is ready for real-world applications, efficient implementation is crucial. We first analyze the time breakdown of representative block-level selective sparse attention method, InfLLM Xiao et al. (2024a). From (1)(2)(3) in Fig. 5, we can observe that although selective sparse attention can significantly reduce the complexity of attention calculations, the actual computation time is still highly dependent on the implementation. The incompatibility with efficient attention implementations such as Flash Attention has resulted in methods requiring historical attention scores Zhang et al. (2024c); Oren et al. (2024); Li et al. (2024); Xiao et al. (2024a) being difficult to be applied in real-world Web applications. Through the analysis of InfLLM’s Flash Attention-compatible version, we make several discoveries. The initial motivation for estimating token criticality at the block-level is to reduce the overhead of selection function (mainly considering dot product calculation). However, we find that dot product is not the primary performance bottleneck. Instead, a significant portion of the overhead comes from indexing the KV Cache using selected indices and making them contiguous in GPU memory, which frequently occurs during the updating of KV blocks and the concatenation of selected KV Cache. The extensive I/O required for this operation further exacerbates the memory-bound in LLMs inference. Based on this, we propose that Paged Attention is a more suitable implementation for selective sparse attention. Using Paged KV Cache management (with page size=1 for TokenSelect), we can reduce the I/O volume for selection results from the scale of all selected KV Caches $O(2kd)$ to the scale of their indices $O(k)$. However, by observing (4) in Fig. 5, we find that we encounter another bottleneck under Paged KV Cache management. Since logically contiguous KV Cache is not entirely contiguous in GPU memory, it also needs to be made contiguous before performing computational operations. To address this issue, we draw inspiration from the concept of Paged Attention and implement a Paged Dot Product Kernel using Triton Tillet et al. (2019), which significantly improves the overall efficiency of TokenSelect.
Figure 5: Time breakdown for single chunk prefill step under different attention implementations (chunk size: 512, KV Cache length: 128K, attended tokens: 4K).
6 Experiments
In this section, we introduce the experimental setup and evaluate the performance and efficiency of our TokenSelect on long-context inference benchmarks.
6.1 Experimental Settings
@@
To evaluate TokenSelect’s performance on long-context inference, we use three representative datasets: InfiniteBench Zhang et al. (2024a), RULER Hsieh et al. (2024), and LongBench Bai et al. (2024). Detailed descriptions and the evaluation metrics used are provided in Appendix C.
@@
To conduct a comprehensive evaluation of TokenSelect’s performance, we carry out benchmarks on three mainstream open-source LLMs - Qwen2-7B-Instruct Yang et al. (2024a), Llama-3-8B-Instruct Grattafiori et al. (2024), and Yi-1.5-6B-Chat AI et al. (2024) - comparing against the following state-of-the-art long-context inference methods: NTK-aware scaled RoPE, SelfExtend, StreamingLLM, InfLLM and MInference. Detailed descriptions of these methods are provided in Appendix B. It is worth noting that since MInference doesn’t support length extrapolation, we use an alternative evaluation method, applying it to Llama-3-8B-Instruct-262k (Llama3 after long-text post-training). Additionally, we do not include another state-of-the-art method, QUEST, as it does not support Grouped Query Attention.
Implementation details.
[…]
6.2 Performance Comparisons
InfiniteBench.
As shown in Table 1, our TokenSelect achieves significantly superior overall performance on InfiniteBench compared to all baseline methods, even though TokenSelect uses the smallest token budget (<3K). The fact that it significantly outperforms the original models demonstrates TokenSelect’s strong length extrapolation capability. We analyze that this is due to our adoption of a fine-grained KV Cache selection strategy, while considering the equal contribution of each head to selection, which ensures that we can select most critical tokens. Observing the performance of other methods, we find that RoPE interpolation methods (NTK, SelfExtend) generally perform poorly unless used on specially trained models such as Qwen2-7B-Instruct. The better performance of Qwen2-7B-Instruct on the original model can also be attributed to this. The sparse attention method StreamingLLM, based on fixed sparse patterns, can guarantee some of the model’s capabilities, but due to discarding a large amount of long-context information, it performs poorly on retrieval-related tasks (R.PK, R.Num, R.KV). The block-level selection method InfLLM can retain more long-context information compared to StreamingLLM. However, due to its sub-optimal block-level selection, it results in lower performance on most tasks compared to TokenSelect, even though we set a larger token budget for InfLLM. It is worth noting that Yi-1.5-6B does not perform normally on the R.KV task, as it is unable to correctly recite strings like the UUID.